Determine the first order partial derivative of the following functions:

1.z=in(x+t²).

2.f(x,y,z)=xy²e^-xz.


1
Expert's answer
2021-06-28T03:31:31-0400

1.


"z=\ln(x+t^2)"


"z_x=\dfrac{\partial z}{\partial x}=\dfrac{1}{x+t^2}"

"z_t=\dfrac{\partial z}{\partial t}=\dfrac{2t}{x+t^2}"



2.


"f(x,y,z)=xy^2e^{-xz}"

"f_x=\dfrac{\partial f}{\partial x}=y^2e^{-xz}-xy^2ze^{-xz}"

"f_y=\dfrac{\partial f}{\partial y}=2xye^{-xz}"

"f_z=\dfrac{\partial f}{\partial x}=-x^2y^2e^{-xz}"